For 40 days in the summer, you are working in a small, student-run company that sends out merchandise with university branding to alumni around the world. Every day, you take a sample of 50 shipments that are ready to be shipped to the alumni and inspect them for correctness. Across all days, the average percentage of incorrect shipments is 5 percent. What would be the center line for a p-chart?

40

50

0.05

2.5

None of the above
Answer #1

Sample(n) \(=50\) shipments

Average incorrect percentage \(=5 \%\)

Defects in sample(np) \(=\frac{5}{100} \times 50\)

Center line of \(\mathrm{p}\)-chart \(=\mathrm{P}\) bar \(=\frac{\sum \mathrm{np}}{\sum \mathrm{n}}\) \(=\frac{5 \times 50}{100 \times 50}\)

\(=0.05\)

Therefore, answer is \(0.05\).